1) If 10.0 mL of glacial acetic acid (pure HC2H3O2) is diluted to 1.50 L with water, what is the pH of the resulting solution? The density of glacial acetic acid is 1.05 g/mL.

2) A 0.190 M solution of a weak acid (HA) has a pH of 2.92. Calculate the acid ionization constant (Ka) for the acid.

3) Determine the percent ionization of a 0.250 M solution of benzoic acid.
1
Expert's answer
2020-04-06T12:06:32-0400

1) glacial acetic acid is 100% acetic acid.

m(CH3COOH) = 10*1.05 = 10.5 g

n(CH3COOH) = "{\frac {10.5} {60}}=0.175" mol

C(CH3COOH) ="{\frac {0.175} {1.5}}" = 0.1167 M

x M of acid had dissociated:

0.1167-x x x

CH3COOH = CH3COO- + H+

if Ka(CH3COOH) = 1.74*10-5 then

Ka = "{\frac {[H^+][CH_3COO^-]} {[CH_3COOH]}}={\frac {x^2} {0.1167-x}}=1.74*10^{-5}"

we neglect "-x" in denominator and resulting equation is

"{\frac {x^2} {0.1167}}=1.74*10^{-5}"

x2 = 1.74*10-5*0.1167 = 2.03*10-6

x =[H+]= 1.42*10-3

pH = -lg[H+] = -lg(1.42*10-3)=2.85

2)

HA = H+ + A-

if x M of HA had dissociated then

[H+] = [A-] = x

[HA] = 0.190 - x

Ka = "{\frac {[H^+][A^-]} {[HA]}}={\frac {[x]^2} {(0.190-x)}}"

[H+] = x = 10-pH = 10-2.92 = 0.0012 M

then Ka = "{\frac {(0.0012)^2} {(0.190-0.0012)}}=7.63*10^{-6}"

3)

Ka(C6H5COOH) = 6.17*10-5

if x M of benzoic acid had dissociated then

0.250 - x x x

C6H5COOH = C6H5COO- + H+

Ka = "{\frac {[H^+][C_6H_5COO^-]} {[C6H5COOH]}}={\frac {x^2} {0.250-x}}=6.17*10^{-5}"

After solving the equation we get

x = 3.9*10-3 M

The percent ionization is "{\frac {3.9*10^{-3}} {0.25}}=1.56%" %